A lead-acid battery uses a redox reaction in which lead(0) and lead(IV) are both converted to lead(II). This reaction is facilitated by the presence of sulfuric acid, H2SO4, as shown by the reaction Pb+PbO2+2H2SO4→2PbSO4+2H2O.

Suppose that a fully charged lead-acid battery contains 1.50 L of 5.00 M H2SO4. What will be the concentration of H2SO4 in the battery after 2.80 A of current is drawn from the battery for 5.50 hours ?
Answer #1

charge = current x time

             = 2.80 x 5.50 x 60 x 60

           = 55440 C

moles of electrons utlised = charge / faraday

                                         = 55440 / 96485

                                         = 0.575

moles of H2SO4 = moles of electrons = 0.575

initial moles = 1.5 x 5 = 7.5

moles remains = 7.5 - 0.575 = 6.925

concentration of H2SO4 = moles / volume

                                     = 6.925 / 1.50

                                    = 4.62 M

concentration of H2SO4 = 4.62 M